OpenStudy (anonymous):

Just a few more! I must get my grade up, only hours left! im timed D: A quadratic: 4/x + 1/x-2 =1

OpenStudy (anonymous):

do you want to learn how to do it or just get an answer?

OpenStudy (anonymous):

u get a good answer for just asking that question ktklown

OpenStudy (anonymous):

i think this is \[\frac{4}{x}+\frac{1}{x-2}=1\] in which case you get a quadratic.

OpenStudy (anonymous):

yes you are :)

OpenStudy (anonymous):

yup

OpenStudy (lilg132):

(puts his head down in shame and goes away for a while)

OpenStudy (anonymous):

@starshine if you just want the answers, type carefully and you can get them easily without asking anyone or getting answers that may be incorrect. use parentheses where needed. see my example here http://www.wolframalpha.com/input/?i=4%2Fx+%2B+1%2F%28x-2%29+%3D1

OpenStudy (lilg132):

sorry for my incorrect answer was just having a go lol

OpenStudy (anonymous):

Yes, if you just want the answers don't bother coming on here, you can get much faster and more accurate answers fro wolfram alpha.

OpenStudy (anonymous):

I would love to know it, but i really just need answers right now, im timed, its vital i get them, i really really apprecieate any help :)

OpenStudy (anonymous):

OK, so if you just need answers, wolfram alpha is better than here. You can always come back here when you want to learn :-)

OpenStudy (anonymous):

Literally just type in your equation and press return and it will solve it for you.

OpenStudy (anonymous):

The URL is: http://www.wolframalpha.com

OpenStudy (anonymous):

it doesnt work

OpenStudy (anonymous):

what doesn't work? wolfram alpha?

OpenStudy (anonymous):

it wont give me the answer i need

OpenStudy (anonymous):

show me your equation and what you typed in. You often need paranthesis around terms.

hero (hero):

After she gets the answers, she will no longer be interested in learning it.. Just a guess

OpenStudy (anonymous):

she'll be no less interested than she is today

OpenStudy (anonymous):

I would rather spend my time on people here who want to learn, like nurse10 (in another window) and shanon123 (in a third window)

hero (hero):

What's a window?

OpenStudy (anonymous):

a rectangular area of a screen on a computer used to interact with a computer program

OpenStudy (anonymous):

I have to learn it!!!! You guys wont be able to help me on a test